# Leave No One Behind: Enhancing Diversity While Maintaining Accuracy in Social Recommendation

### v1

@@ Abstract
Social recommendation, a branch of algorithms that utilizes social connection information to construct recommender systems, has demonstrated its effectiveness in enhancing recommendation accuracy. However, apart from accuracy, the diversity of recommendations also plays a critical role in user engagement. Unfortunately, the impact of social recommendation models on recommendation diversity remains largely unexplored. In this study, we investigate the dual performance of existing social recommendation algorithms in terms of accuracy and diversity. Our empirical findings highlight a concerning trend: social recommendation models tend to decrease diversity, despite their accuracy improvements. To address this issue, we propose a novel approach called Diversified Social Recommendation (DivSR), which leverages relational knowledge distillation techniques to transfer high-diversity structured knowledge from non-social recommendation models to social recommendation models. DivSR is designed as a simple, model-agnostic framework that integrates seamlessly with existing social recommendation architectures. Experimental results on three benchmark datasets demonstrate that DivSR significantly increases diversity without markedly compromising accuracy across various social recommendation backbones, achieving a better accuracy-diversity trade-off. Our code and data are publicly available at: https://github.com/ll0ruc/DivSR.
@@
Recommender Systems Social Recommendation Diversified Recommendation
@@ 1 Introduction
In this era of information explosion, recommender systems play a vital role in alleviating the problem of information overload for users [7, 42]. Recommender systems have been successfully implemented across various domains including e-commerce [29], online news [37], and multimedia contents [23]. With the advancement of recommendation algorithms, accuracy serves as the dominant target or even the only target to maximize the utility of recommendation systems [1, 12]. As one of its typical representatives, social recommendation [18, 31] utilizes social resources such as interpersonal relations and influence, as extra information to enhance performance. This approach typically incorporates social connections either as a regularization term [11] or through feature embeddings from neighboring nodes using a graph attention network framework [8].
@@
However, an accurate recommendation is not necessarily a satisfactory one [2]. Users on e-commerce platforms seek more than just precisely accurate products; they also devote significant time to exploring news-feed products for diverse options. Thus, an ideal recommendation system should fulfill both accuracy and diversity requirements [22]. Unfortunately, existing social recommendation systems primarily focus on enhancing accuracy, often overlooking diversity. This oversight is problematic considering that social recommendation systems are fundamentally influenced by social influence theory [21], which posits that users influenced by their social connections tend to adopt similar preferences. This can lead to homogeneity in the recommendations over time due to the over-reliance on preferences within a user’s immediate social circle [28], underlining the necessity of fostering diversity in social recommendations.
@@
In this work, we revisit the current social recommendation algorithms, probing their performance in terms of diversity. We conducted extensive experiments to evaluate the performance of several existing social recommendation methods on three benchmark datasets in terms of both accuracy and diversity. For each algorithm, we removed socially-relevant modules, enabling it to make item recommendations without utilizing social information. Such a comparison between social recommendation methods and their variant (non-social methods) allows us to discern performance discrepancies attributable to the integration of social relationships. Our empirical findings reveal that existing social recommendation models tend to decrease diversity while improving accuracy compared to non-social recommendation methods.
@@
To achieve a better accuracy-diversity trade-off on social recommendation, we propose the DivSR (Diversified Social Recommendation) framework, which leverages knowledge distillation to achieve system-level overall diversity in recommendations. Fundamentally, DivSR maintains a model-agnostic design, allowing seamless integration with various social recommendation backbone models. In DivSR, we train a social recommendation model as the student model to combine high accuracy and high diversity. The high diversity is derived from a pre-trained teacher model, which is a non-social recommendation counterpart. We design a knowledge transfer module using relational distillation learning technology [24], which distills structured similarity knowledge between users and friends from the teacher model to the student model. To achieve a balanced trade-off between accuracy and diversity, DivSR optimizes both the recommendation task and the knowledge distillation task simultaneously within a primary and auxiliary learning framework.
@@
We perform experiments on three widely used public datasets, incorporating five robust social recommenders as backbone models. Comprehensive results demonstrate that DivSR enhances diversity without excessively sacrificing accuracy across various social recommendation backbones. DivSR demonstrates a superior accuracy-diversity trade-off compared to several diversified models. The key contributions of this paper are summarized as follows:
@@
  • We empirically assess the accuracy-diversity performance of social recommendation systems, finding that they typically reduce diversity compared to their non-social counterparts.
@@
  • We propose DivSR, a model-agnostic framework utilizing knowledge distillation techniques designed to foster diversity in social recommendations. This approach includes a diversity-knowledge transfer module capable of distilling structured similarity information.
@@
  • Through experiments on three datasets, we demonstrate the effectiveness of DivSR, highlighting its ability to significantly enhance diversity without excessively sacrificing accuracy across various social recommendation systems.
@@ 2.1 Social Recommendation
To improve the accuracy of recommendation results, numerous social recommendation methods have been developed, incorporating online social relationships between users as side information [9, 16]. Early models like SoReg [19] and SocialMF [11] integrated social connections as regularization terms or utilized trust relationships to project users into latent representations, exemplified by TrustMF [36] and TrustSVD [9]. In recent years, graph neural networks (GNNs) [15] have achieved great success in deep learning, owing to their strong capability on modeling graph data. DiffNet [35] and its extension DiffNet++ [34] modeled the information diffusion process in social graph to enlarge the users’ influence scope. Multi-channel hypergraph convolutional network was employed on MHCN [41] to enhance social recommendation by leveraging high-order user relations. SEPT [40] proposed a socially-aware self-supervised learning(SSL) framework that discovers supervisory signals from two complementary views of raw data. DESIGN [31] introduced knowledge distillation between models that rely on different data sources to leverage social information effectively. Wang et al. [32] proposed a universal denoised self-augmented learning framework that incorporates social influence to decipher user preferences while mitigating noisy effects. Nevertheless, these methods mostly aim to improve accuracy while neglecting diversity. Our work contributes to achieving a balance between accuracy and diversity in social recommendation.
@@ 2.2 Diversified Recommendation
Diversified recommendation aims to provide users with a more varied set of items, enabling users to discover new and unexplored interests [6, 17]. The accuracy-diversity dilemma, pointing higher accuracy often means losing diversity to some extent and vice versa. A classical re-ranking work to enhance diversity is maximal marginal relevance (MMR) [4], which used the notion of marginal relevance to combine relevance and diversity with a trade-off parameter. Determinantal point process (DPP) [5] re-ranked items to achieve the largest determinant on the item’s similarity matrix. Zheng et al. [43] introduced an end-to-end diversified recommendation model named DGCN, which selects node neighbors based on the inverse category frequency for diverse aggregation and further utilizes category-boosted negative sampling and adversarial learning to diverse items in the embedding space. DGRec [38] targeted diversifying GNN-based recommender systems with diversified embedding generation. Different from these methods, our work facilitates seamless integration with various social recommendation systems, effectively enhancing accuracy while maintaining high diversity when integrating social relationships.
@@ 3.1 Problem Statement
In the task of social recommendation, we denote $\textit{U}=\{u_{1},u_{2},\ldots,u_{M}\}$ (|U| = M) as the user set and $\textit{V}=\{v_{1},v_{2},\ldots,v_{N}\}$ (|V| = N) as the item set. We use $a$, $b$ to index users, and $i$, $j$ to index items. Let $G_{s}=<U,S>$ denote a directed social graph, where $S\in\mathbb{R}^{M\times M}$ is a matrix representing social relations between users. For each user-user pair ($a,b$), $s_{ab}$ = 1 if user $a$ trusts user $b$ and 0 otherwise. Let $G_{r}=<U\cup V,R>$ denote an undirected bipartite graph, where $R\in\mathbb{R}^{M\times N}$ is a user-item interaction matrix. For each user-item pair ($a,i$), $r_{ai}$ = 1 indicates that user $a$ consumes item $i$ and 0 otherwise. The core objective of a social recommendation system leveraging $G_{r}$ and $G_{s}$ is to predict and recommend top $k$ interested items that a user is likely to be interested in, based on their past interactions and social influences. The diversified social recommendation task aims to recommend items that users prefer while ensuring high system-level overall diversity.
@@ 3.2 Accuracy-Diversity Dilemma in Social Recommendation
In this paper, we revisit the current social recommendation algorithms, probing their performance in terms of diversity. Nevertheless, the model performance comparison between different social recommendation models is not our primary interest. Instead, we are particularly concerned with evaluating how accuracy and diversity are influenced by the introduction of social relationships. We conduct preliminary experiments to analyze the two-fold performance w.r.t. accuracy (Recall@100) and diversity (Coverage@100) on three widely-used public datasets (Yelp [35], Ciao [30], and Flickr [33]) between Social Recommender System (Social RS) and Non-social Recommender System (Non-social RS).
@@
For Social RS, we choose several typical social recommendation models, including TrustMF [36], SocialMF [11], DiffNet [35], DESIGN [31]), and MHCN [41]. In contrast, we just remove the socially relevant modules from each Social RS while keeping all other components unchanged to get a non-social recommendation counterpart. For example, DiffNet utilizes a layer-wise GNN structure to simulate the recursive social diffusion process. The final user embedding $p_{a}$ is composed of two parts: the embedding from social diffusion layers and the preferences from historical behaviors:
@@
where $h_{a}^{0}$ and $q_{i}$ represent the initial free embedding of user $a$ and item $i$, respectively. $\mathcal{N}_{r}(a)$ is the itemset that user $a$ consumed, $GNN(\cdot)$ is a layer-wise graph neural network and $K$ indicates the number of GCN layers. When the social module is removed, $h_{a}^{K}$ simply defaults to $h_{a}^{0}$, as the social diffusion layers that would normally utilize the explicit user-user social graph ($G_{s}$) are omitted. Thus, DiffNet (w/o social) relies solely on the user-item interaction graph ($G_{r}$) for generating recommendations.
From Table 1, we observe that social recommendation models can significantly improve the accuracy compared to their non-social variants. For example, SocialMF shows an improvement of approximately 12% on Yelp and 5% on Ciao when compared to SocialMF (w/o social). In terms of diversity, most social recommendation methods severely reduce system aggregate-level diversity. For instance, the diversity of DiffNet drops from 40% to 17% on Yelp and from 39% to 28% on Ciao when compared to its non-social variant. From these findings, it can be concluded that: Social recommendation methods usually reduce recommendation diversity while improving recommendation accuracy compared to their non-social variants.
@@ 3.3 Embedding Similarity in Social Recommendation
To gain more insights into the findings obtained in Table 1 and understand why SocialRS reduces diversity, we analyze embedding similarities between users and their friends as training progresses. Following [28], we calculate the cosine similarity between the feature vectors of two users, normalized to $0-1$. A higher value indicates greater similarity.
Figure 1 illustrates changes in embedding similarity, accuracy, and diversity during training DiffNet on Yelp dataset. Initially, with random vector initialization, similarity is low and diversity is high. As training advances, vector similarities increase, boosting accuracy but lowering diversity. In advanced training stages, vector similarities peak and then decline, whereas diversity decreases to a minimum before starting to increase again. Accuracy, meanwhile, converges towards a stable value. Extended training post-convergence may lead to decreased accuracy, with embedding similarities falling and diversity rising. This pattern suggests a negative correlation between embedding similarity and diversity. Furthermore, we also observe that social recommendation methods always generate more similar representations (higher embedding similarity) between users and their friends compared to their non-social variants.
@@ 4.1 Overall Framework
In this paper, we propose a Diversified Social Recommendation framework based on knowledge distillation, (named DivSR), which achieves a more balanced trade-off between accuracy and diversity. The overall architecture is shown in Figure 2. DivSR is motivated by two advantages: i) Social RS typically exhibits satisfactory recommendation accuracy, and ii) Non-social RS tends to offer higher recommendation diversity. To capitalize on these strengths, DivSR adopts a novel approach wherein a social recommendation model serves as the student model, amalgamating the strengths of both high accuracy and high diversity. The high diversity is derived from a pre-trained teacher model, which is a non-social recommendation counterpart. We design a knowledge transfer module based on relational distillation learning technology, which facilitates the transfer of structured diversity knowledge from the teacher model to the student model. Specifically, the structured vector similarity between users and friends is utilized to characterize diversity knowledge, as revealed in Section 3.3, where a lower vector similarity corresponds to an increase in diversity. DivSR optimizes both the recommendation task and the knowledge distillation task simultaneously within a primary and auxiliary learning framework.
Figure 2: The left is an overview of the proposed DivSR framework. The right is the Diversity-Knowledge Transfer module.
@@ 4.2 Teacher Model & Student Model
DivSR is a simple and model-agnostic solution that can be easily deployed on existing social recommendation models. Given a social recommendation method, we first pre-train its corresponding non-social variant as the teacher network under the recommendation loss. Then we jointly train the social recommendation method as the student model using both recommendation loss and distillation loss. During the training phase of the student network, the teacher network has already been fully trained and frozen.
Specifically, we formalize $f_{t}(\theta_{t})$ as a Non-social RS, where $\theta_{t}$ is the model parameters of the teacher network. Since the teacher model does not consider social relationships, we provide it with a user-item bipartite graph. The teacher model is trained under the Bayesian Personalized Ranking (BPR) [26] loss:
@@
where $\mathcal{N}_{r}(a)$ denotes the itemsets user $a$ consumed, $\sigma(\cdot)$ is a sigmoid function and $\lambda$ is a regularization parameter that prevents overfitting. The rating score $\tilde{r}_{a,i}(\theta_{t})=(q_{i}^{t})^{T}p_{a}^{t}$ is determined by the final user/item embedding ($p_{a}^{t}$, $q_{i}^{t}$) produced by the teacher model.
Now, considering a student model $f_{s}(\theta_{s})$, which incorporates both the user-item bipartite graph ($G_{r}$) and the user-user social graph ($G_{s}$), we proceed with joint training on both the recommendation task and the knowledge distillation task. The knowledge distillation task will be introduced later and the objective function for the recommendation task is same as Equation 3. The final user/item embedding generated by student model for user $a$, item $i$ are $p_{a}^{s}$, $q_{i}^{s}$, whose dot product forms the rating scoring, $\tilde{r}_{a,i}(\theta_{s})=(q_{i}^{s})^{T}p_{a}^{s}$.

### v2

@@ Abstract
Social recommendation, which incorporates social connections into recommender systems, has proven effective in improving recommendation accuracy. However, beyond accuracy, diversity is also crucial for enhancing user engagement. Despite its importance, the impact of social recommendation models on diversity remains largely unexplored. In this study, we systematically examine the dual performance of existing social recommendation algorithms in terms of both accuracy and diversity. Our empirical analysis reveals a concerning trend: while social recommendation models enhance accuracy, they often reduce diversity. To address this issue, we propose Diversified Social Recommendation (DivSR), a novel approach that employs relational knowledge distillation to transfer high-diversity structured knowledge from non-social recommendation models to social recommendation models. DivSR is a lightweight, model-agnostic framework that seamlessly integrates with existing social recommendation architectures. Experiments on three benchmark datasets demonstrate that DivSR significantly enhances diversity while maintaining competitive accuracy, achieving a superior accuracy-diversity trade-off. Our code and data are publicly available at: https://github.com/ll0ruc/DivSR.
@@
Recommender Systems Social Recommendation Recommendation Diversity
@@ 1 Introduction
In the era of information overload, recommender systems play a crucial role in helping users navigate vast amounts of content [7]. They have been successfully deployed across various domains, including e-commerce [29], online news [37], and multimedia content [23]. With the advancement of recommendation algorithms, accuracy has become the primary, and often sole, optimization objective [1, 12]. As a major branch of recommender systems, social recommendation [18, 31] leverages social resources, such as interpersonal relationships and influence, to enhance recommendation performance. This approach typically incorporates social connections either as a regularization constraint [11] or by extracting feature embeddings from neighboring nodes using a graph attention network framework [8].
@@
However, an accurate recommendation is not necessarily a satisfactory one [2]. Users on e-commerce platforms seek more than just highly relevant products; they also devote significant time to exploring news-feed products for a wider range of options. Thus, an ideal recommendation system should fulfill both accuracy and diversity requirements [22]. Unfortunately, existing social recommendation systems primarily focus on enhancing accuracy, often overlooking diversity. This oversight is problematic considering that social recommendation systems are fundamentally influenced by social influence theory [21], which posits that users influenced by their social connections tend to adopt similar preferences. This can lead to homogeneity in the recommendations over time due to the over-reliance on preferences within a user’s immediate social circle [28], underlining the necessity of fostering diversity in social recommendations.
@@
In this work, we revisit current social recommendation algorithms, probing their performance in terms of diversity. We conduct extensive experiments to evaluate the performance of several existing social recommendation methods on three benchmark datasets assessing both accuracy and diversity. For each algorithm, we remove socially-relevant modules, enabling it to make item recommendations without utilizing social information. By comparing social recommendation methods with their non-social variants, we can identify performance discrepancies resulting from the integration of social relationships. Our empirical findings reveal that existing social recommendation models tend to decrease diversity while improving accuracy compared to non-social recommendation methods.
@@
To achieve a better accuracy-diversity trade-off in social recommendation, we propose the DivSR (Diversified Social Recommendation) framework, which leverages knowledge distillation to achieve system-level overall diversity in recommendations. Fundamentally, DivSR maintains a model-agnostic design, allowing seamless integration with various social recommendation backbone models. In DivSR, we train a social recommendation model as the student model to combine high accuracy and high diversity. The high diversity is derived from a pre-trained teacher model, which serves as a non-social recommendation counterpart. We design a knowledge transfer module using relational distillation learning technology [24], which distills structured similarity knowledge between users and their social connections from the teacher model to the student model. To strike a balanced trade-off between accuracy and diversity, DivSR optimizes both the recommendation task and the knowledge distillation task simultaneously within a primary and auxiliary learning framework.
@@
We perform experiments on three widely used public datasets, incorporating five robust social recommenders as backbone models. Comprehensive results demonstrate that DivSR enhances diversity without significantly sacrificing accuracy across various social recommendation backbones. DivSR achieves a superior accuracy-diversity trade-off compared to several diversified models. The key contributions of this paper are summarized as follows:
@@
  • We empirically assess the accuracy-diversity performance of social recommendation systems and find that they typically reduce diversity compared to their non-social counterparts.
@@
  • We propose DivSR, a model-agnostic framework that leverages knowledge distillation techniques to foster diversity in social recommendations. This approach includes a diversity-knowledge transfer module that distills structured similarity information.
@@
  • Through experiments on three datasets, we demonstrate the effectiveness of DivSR, highlighting its ability to significantly enhance diversity without significantly sacrificing accuracy across various social recommendation systems.
@@ 2.1 Social Recommendation
To improve the accuracy of recommendation results, numerous social recommendation methods have been developed, incorporating online social relationships between users as side information [9, 16]. Early models like SoReg [19] and SocialMF [11] integrate social connections as regularization terms or utilize trust relationships to project users into latent representations. In recent years, graph neural networks (GNNs) [15] have achieved great success in deep learning, owing to their strong capability on modeling graph data. DiffNet [35] and its extension DiffNet++ [34] model the information diffusion process in social graphs to enlarge the users’ influence scope. Multi-channel hypergraph convolutional network is employed on MHCN [41] to enhance social recommendation by leveraging high-order user relations. Self-supervised learning (SSL) is utilized in SEPT [40] to improve social recommendation by uncovering supervisory signals from two complementary views of raw data. DESIGN [31] introduces knowledge distillation between models that rely on different data sources to leverage social information effectively. A universal denoised self-augmented learning framework (DSL) [32] incorporates social influence to decipher user preferences while mitigating noisy effects. Nevertheless, these methods mostly aim to improve accuracy while neglecting diversity. Our work contributes to achieving a balance between accuracy and diversity in social recommendation.
@@ 2.2 Diversified Recommendation
Diversified recommendation aims to provide users with a more varied set of items, enabling users to discover new and unexplored interests [6, 17]. The accuracy-diversity dilemma, pointing higher accuracy often means losing diversity to some extent and vice versa. A classical re-ranking work to enhance diversity is maximal marginal relevance (MMR) [4], which uses the notion of marginal relevance to combine relevance and diversity with a trade-off parameter. Determinantal point process (DPP) [5] re-ranks items to achieve the largest determinant on the item’s similarity matrix. DGCN [42] selects node neighbors based on the inverse category frequency for diverse aggregation and further utilizes category-boosted negative sampling and adversarial learning to diverse items in the embedding space. DGRec [38] targets diversifying GNN-based recommender systems with diversified embedding generation. Different from these methods, our work facilitates seamless integration with various social recommendation systems, effectively enhancing accuracy while maintaining high diversity.
@@ 3.1 Problem Statement
In the task of social recommendation, let $\textit{U}=\{u_{1},u_{2},\ldots,u_{M}\}$ (|U| = M) represent the set of users, and $\textit{V}=\{v_{1},v_{2},\ldots,v_{N}\}$ (|V| = N) represent the set of items. We use indices $a$ and $b$ to refer to users, and $i$ and $j$ to refer to items. Let $G_{s}=<U,S>$ denote a directed social graph, where $S\in\mathbb{R}^{M\times M}$ is a matrix representing social relations between users. For each user-user pair ($a,b$), $s_{ab}$ = 1 if user $a$ trusts user $b$ and 0 otherwise. Let $G_{r}=<U\cup V,R>$ denote an undirected bipartite graph, where $R\in\mathbb{R}^{M\times N}$ is a user-item interaction matrix. For each user-item pair ($a,i$), $r_{ai}$ = 1 indicates that user $a$ has interacted with item $i$ and 0 otherwise. The primary objective of a social recommendation system leveraging $G_{r}$ and $G_{s}$ is to predict and recommend the top $k$ items a user is likely to be interested in, based on their past interactions and social influences. The diversified social recommendation task aims to recommend items that users prefer while ensuring high system-level overall diversity.
@@ 3.2 Accuracy-Diversity Dilemma in Social Recommendation
In this paper, we revisit existing social recommendation algorithms, probing their performance in terms of accuracy and diversity. Nevertheless, the comparison of model performance across different social recommendation approaches is not the primary focus of this study. Instead, we specifically examine how the incorporation of social relationships influences both accuracy and diversity. We conduct preliminary experiments to evaluate the two-fold performance w.r.t. accuracy (Recall@100) and diversity (Coverage@100) between Social Recommender System (Social RS) and Non-social Recommender System (Non-social RS). These experiments are conducted on three widely-used public datasets (Yelp [35], Ciao [30], and Flickr [33]).
@@
For Social RS, we select several representative social recommendation models, including TrustMF [36], SocialMF [11], DiffNet [35], DESIGN [31], and MHCN [41]. In contrast, to create the non-social recommendation counterpart, we remove the socially relevant modules from each Social RS while keeping all other components intact. For instance, DiffNet utilizes a layer-wise GNN structure to simulate the recursive social diffusion process. The final user embedding $p_{a}$ consists of two parts: the embedding derived from social diffusion layers and the preferences based on historical interactions:
@@
where $h_{a}^{0}$ and $q_{i}$ represent the initial free embedding of user $a$ and item $i$, respectively. $\mathcal{N}_{r}(a)$ is the itemset that user $a$ consumed, $GNN(\cdot)$ denotes a layer-wise graph neural network, and $K$ indicates the number of GCN layers. When the social module is removed, $h_{a}^{K}$ simply defaults to $h_{a}^{0}$, as the social diffusion layers, which would normally leverage the explicit user-user social graph ($G_{s}$), are omitted. Therefore, DiffNet (w/o social) relies exclusively on the user-item interaction graph ($G_{r}$) to generate recommendations.
From Table 1, we observe that social recommendation models significantly enhance accuracy compared to their non-social counterparts. For instance, SocialMF shows an improvement of approximately 12% on Yelp and 5% on Ciao when compared to SocialMF (w/o social). However, in terms of diversity, most social recommendation methods lead to a substantial decrease in system-level diversity. For example, the diversity of DiffNet drops from 40% to 17% on Yelp and from 39% to 28% on Ciao when compared to its non-social variant. Based on these results, we can conclude that: social recommendation methods usually reduce recommendation diversity while improving recommendation accuracy compared to their non-social variants.
@@ 3.3 Embedding Similarity in Social Recommendation
To further explore the findings presented in Table 1 and understand why SocialRS reduces diversity, we analyze the embedding similarities between users and their friends. Following [28], we calculate the cosine similarity between the feature vectors of two users, normalizing the values to a range of 0 to 1, where higher values indicate greater similarity.
Figure 1 shows the changes in embedding similarity, accuracy, and diversity during the training of DiffNet on Yelp dataset. Initially, with random vector initialization, similarity is low, and diversity is high. As training progresses, vector similarities increase, leading to improved accuracy but reduced diversity. In the later stages of training, embedding similarities reach their peak and then begin to decline, while diversity drops to a minimum before gradually increasing again. Meanwhile, accuracy converges toward a stable value. If training continues past convergence, accuracy may decrease as embedding similarities drop and diversity increases. This pattern suggests a negative correlation between embedding similarity and diversity. Additionally, we also observe that social recommendation methods always generate more similar embeddings (higher embedding similarity) between users and their friends compared to their non-social variants.
@@ 4.1 Overall Framework
In this paper, we propose a Diversified Social Recommendation framework based on knowledge distillation, named DivSR, which achieves a more balanced trade-off between recommendation accuracy and diversity. The overall architecture is shown in Figure 2. DivSR is motivated by two key advantages: i) Social RS typically exhibits satisfactory recommendation accuracy, and ii) Non-social RS tends to offer higher recommendation diversity. To leverage these strengths, DivSR introduces a novel approach where a social recommendation model serves as the student model, combining the benefits of both high accuracy and high diversity. The high diversity is derived from a pre-trained teacher model, which is its non-social recommendation counterpart. We design a knowledge transfer module based on relational distillation learning, which facilitates the transfer of structured diversity knowledge from the teacher model to the student model. Specifically, the structured vector similarity between users and their friends is used to capture diversity knowledge, as described in Section 3.3, where lower vector similarity corresponds to higher diversity. DivSR optimizes both the recommendation task and the knowledge distillation task simultaneously within a primary and auxiliary learning framework.
Figure 2: The overall framework of DivSR.
@@ 4.2 Teacher Model & Student Model
DivSR is a simple and model-agnostic solution that can be easily deployed on existing social recommendation models. Given a social recommendation method, we first pre-train its corresponding non-social variant as the teacher model. using the recommendation loss. Then, we jointly train the social recommendation model as the student model, using both recommendation loss and distillation loss. During the training of the student network, the teacher network has already been fully trained and frozen.
Specifically, we formalize $f_{t}(\theta_{t})$ as a Non-social RS, where $\theta_{t}$ is the model parameters of the teacher network. Since the teacher model does not account for social relationships, it is provided with a user-item bipartite graph. The teacher model is trained using the Bayesian Personalized Ranking (BPR) [26] loss:
@@
where $\mathcal{N}_{r}(a)$ denotes the set of items consumed by user $a$, $\sigma(\cdot)$ is a sigmoid function, and $\lambda$ is a regularization parameter to prevent overfitting. The rating score $\tilde{r}_{a,i}(\theta_{t})=(q_{i}^{t})^{T}p_{a}^{t}$ is determined by the final user embedding $p_{a}^{t}$ and item embedding $q_{i}^{t}$ produced by the teacher model.
Now, considering the student model $f_{s}(\theta_{s})$, which incorporates both the user-item bipartite graph ($G_{r}$) and the user-user social graph ($G_{s}$), we proceed with joint training on both the recommendation task and the knowledge distillation task. The knowledge distillation task will be introduced later and the objective function for the recommendation task is same as Equation 3. The final user/item embeddings generated by the student model for user $a$ and item $i$ are $p_{a}^{s}$ and $q_{i}^{s}$, respectively, and their dot product forms the rating scoring, $\tilde{r}_{a,i}(\theta_{s})=(q_{i}^{s})^{T}p_{a}^{s}$.